Scenario: Kaleidoscope Kolors, Inc. Kaleidoscope Kolors, Inc., is a San Francisco company that sells clothing all across the U.s. Kaleidoscope has subsidiaries in several nations internationally. At the end of the year Kaleidoscope integrates its financial data from international operations with its headquarters in San Francisco. Suppose Nation A has a currency called the Punto (P). At the start of the year, the exchange rate between the Punto and the U.S. dollar was P150/$. The inflation rate in Nation A is running at an annual rate of 250 percent, whereas inflation in the U.S. is running at 2 percent. Which of the following would most likely be the new exchange rate that Kaleidoscope can expect at the end of the year? Select one: a. P525/$ b. P43.71/$ c. P514.70/$ d. $43.71/P
Answer #1

New exchange rate​​​punto/$ = initial exchange rate​​​punto/$ * [(1+π​​​​​​A​​​)/(1+π​​​​​​​​us)] = 150*[(1+250%)/(1+2%)] = 150*[(1+2.5)/(1+0.02)] = 150*(3.5/1.02) = P514.70/$

Answer: option C